OpenStudy (anonymous):

Let f(x) = 2^k + 9, where k is a real number. If f(3) : f(6) = 1:3, determine the value of f(9) - f(3).

OpenStudy (anonymous):

Tougher problem...watch

OpenStudy (anonymous):

I assume that f(x) = 2^x + 9, not k true?

OpenStudy (agent0smith):

Yeah i'm not sure how we're supposed to do much with this, it seems like k is a constant not a variable...

OpenStudy (anonymous):

so the function f(x) is a constant function?

OpenStudy (agent0smith):

I've no idea, it seems like something is missing though

OpenStudy (anonymous):

That really cant be, becuase if it is a constant function, then f(anything) = to that same constant. It seems otherwise here.

OpenStudy (anonymous):

Then you wouldnt have different ratios..for the same constant, everything will be in a 1:1 ratio

OpenStudy (agent0smith):

Yeah i think there's an x missing somewhere...

OpenStudy (anonymous):

Oh sorry its 2^kx :)

OpenStudy (anonymous):

is it \[2^{k} or k ^{2}\]

OpenStudy (anonymous):

\[2^{kx} + 9\]

OpenStudy (anonymous):

check where the function is inc or dec, from there you can get your answer...

OpenStudy (anonymous):

Let f(x) = \[2^{kx} + 9\], where k is a real number. If f(3) : f(6) = 1:3, determine the value of f(9) - f(3).

OpenStudy (anonymous):

@divu.mkr That's what was only stated here :(

OpenStudy (anonymous):

first, you have to find the value of k. or at least related to k. let's sub the conditions into the equation.\[f(3)=2^{3k}+9\]\[f(6)=2^{6k}+9\] f(3):f(6)=1:3, which means \[3*(2^{3k}+9)=2^{6k}+9\] notice that \[2^{6k}=(2^{3k})^{2}\] now, if we set \[2^{3k}=y\], then we have another equation\[3*(y+9)=y^{2}+9\] solve this for y, we will have the value of \[2^{3k}\] then the rest is easy , just sub in the values, solve the problem.

OpenStudy (anonymous):

didn't do the calculation, i'm gonna do it right now and let you know the answer.

OpenStudy (debbieg):

wait, scratch what I said....

OpenStudy (debbieg):

Yeah, that's what I was trying to do... lol....

OpenStudy (anonymous):

for the equation \[3*(y+9)=y^{2}+9\], you will have y = 6 or -3. since k is a real number, so it's not possible to have \[2^{3k}=-3\]. so we have \[2^{3k}=6\] now we are going to find the value of f(9)-f(3), which is \[2^{9k}+9-(2^{3k}+9)\], which is simplified into \[(2^{3k})^{3}-2^{3k}\] we already know that 2^3k is 6. so the answer is 6^3-6=216-6=210

OpenStudy (debbieg):

Cool problem. Well done, @linkinage .

OpenStudy (anonymous):

@raechelvictoria getting this? if you have questions feel free to ask

OpenStudy (anonymous):

@DebbieG thanks.

OpenStudy (anonymous):

@linkinage Why can't it be (-3)^3 - 6?

OpenStudy (anonymous):

like i explained, since k is a real number, it is not possible having 2^(3k) = -3

OpenStudy (anonymous):

so 2^(3k) = 6 is rock solid and the only case we have to deal with.

OpenStudy (anonymous):

@linkinage Oh okaaaaay. Thank you!

OpenStudy (anonymous):

no worries.